OpenStudy (firejay5):

While at Magic Springs, Katie goes down a frictionless water slide that is 9.0 m above the ground. At the bottom of the slide, Katie reaches a final speed of 13 m/s. What was Katie's initial speed when she pushed off the top of the slide?

OpenStudy (loser66):

I am sorry, I don't think I know how to solve it. To me, the given condition is not enough to solve.

OpenStudy (anonymous):

Hey, can you draw us a picture?

OpenStudy (anonymous):

I would conservation of energy. Define the potential and kinetic energy at the top of the slide and at the bottom. Assuming ideal conditions, the total energy should be conserved. Recall: Potential Energy = m * g * h Kinetic Energy = (1/2) * m * v^2

OpenStudy (anonymous):

Yes, conservation of energy would be also easiest in this case. Look at your initial and final conditions and see what you can cancel out, I also want you to draw a picture because it will show that you understand what's going in the problem.

OpenStudy (perl):

potential energy + kinetic energy is constant

OpenStudy (perl):

PE + KE = k , where k means just constant. where PE = m*g*h KE = 1/2*m*v^2 at the bottom of the slide we have m*g*0 + 1/2*m*13^2 = 169/2 * m Therefor at the top of the slide we must have (let g= 9,81) m*(9.81)*9 + 1/2*m*v^2 = 169/2 *m m ( 9.81 * 9 + 1/2 * v^2 ) = 169/2 * m

OpenStudy (perl):

can you solve for v now?

OpenStudy (perl):

Here is an easier way to look at it Initial kinetic energy + initial potential energy = final kinetic energy + final potential energy Ki + Pi = Kf + Pf 1/2*m*vi^2 + m*g*hi = 1/2mvf^2 + m*g*hf 1/2 *m*vi^2 + m*9.81*9 = 1/2*m(13)^2 + m*g*0

OpenStudy (perl):

the odd thing about this problem, there is no real solution. its possible whoever made this problem did not work out the math